T1ASM(1) T1ASM(1) NAME t1asm−assembletype-1font SYNOPSIS t1asm[−a|−b][−llength][input[output]] DESCRIPTION t1asmassemblesAdobeType1fontprogramsintoeitherPFA(hexadecimal)orPFB(binary) formatsfromahuman-readableform.Ifthefileoutputisnotspecifiedoutputgoestothestandard output.Ifthefileinputisnotspecifiedinputcomesfromthestandardinput. t1asmtokenizesthecharstringdataandperformseexecandcharstringencryptionasspecifiedin the‘‘blackbook,’’AdobeType1FontFormat. Theinputmusthavealineoftheform /-|{stringcurrentfileexchreadstringpop}executeonlydef whichdefinesthecommand,inthiscase‘−|’,thatistostartcharstringdata.Itisanerrornotto definesuchacommand.Anothercommonnameforthiscommandis‘RD’. AfterthestartoftheSubrsarrayintheinput,allopenbraces‘{’notinacommentbegina charstring.Suchacharstringisterminatedbythenextnon-commentclosebrace‘}’.Withinsuch acharstring,onlycomments,integers,andvalidcharstringcommandsareallowed.Valid charstringcommandnamescanbefoundinAdobeType1FontFormatandotherdocuments describingthenewerType2opcodes.Theformatwithinacharstringisunimportantaslongas integersandcommandsareseparatedbyatleastaonewhitespace(space,tab,newline)character. Notethatwithincharstrings,commentsarediscardedbecausetheycannotbeencoded. OPTIONS −−pfa,−a OutputinPFA(ASCII)format. −−pfb,−b OutputinPFB(binary)format.Thisisthedefault. −−block−length=num,−lnum PFBonly:Setthemaximumoutputblocklengthtonum.Thedefaultlengthisaslargeas memoryallows. −−line−length=num,−lnum PFAonly:Setthemaximumlengthofencryptedlinesintheoutputtonum.(Thesearethe linesconsistingwhollyofhexadecimaldigits.)Thedefaultis64. EXAMPLES %t1asmUtopia-Regular.raw>Utopia-Regular.pfb %t1asm-aUtopia-Regular.raw>Utopia-Regular.pfa SEEALSO t1disasm(1),t1ascii(1),t1binary(1),t1unmac(1),t1mac(1) AdobeType1FontFormatisavailablefreefromAdobeasaPDFfile.http://partners.adobe.com/asn/developer/PDFS/TN/T1SPEC.PDF TheType2CharstringFormat,alsoavailablefromAdobeasaPDFfile,describesthenewerType 2operators,whicharealsousedinsomemultiple-masterType1fontslikeAdobeJensonand Kepler.http://partners.adobe.com/asn/developer/PDFS/TN/5177.Type2.pdf AUTHORS LeeHetherington(ilh@lcs.mit.edu) EddieKohler(eddietwo@lcs.mit.edu) 5Mar1998 1